OpenStudy (anonymous):

Question 1 Evaluate 3^3 Question 2 Evaluate 10^1 Question 3 Evaluate -3^2 Question 4 Evaluate 5^0 Question 5 Evaluate 4^3 Question 6 Evaluate -1^1 Question 7 Evaluate -2^6 Question 8 Evaluate 4^3 - 10 Question 9 Evaluate 3^2 + 5 • 3

OpenStudy (anonymous):

1. 27

OpenStudy (anonymous):

2. 10

OpenStudy (anonymous):

27 10 9 1 64

OpenStudy (anonymous):

3. -9

OpenStudy (anonymous):

9) 3^2 = 3 . 3 + 5 • 3 =?

OpenStudy (anonymous):

4. 1

OpenStudy (anonymous):

8) 4 . 4 . 4 - 10 =?

OpenStudy (anonymous):

5. 64

OpenStudy (anonymous):

7) -2 . -2 . -2 . -2 . -2 . -2 =?

OpenStudy (anonymous):

6) -1 . -1 =?

OpenStudy (anonymous):

6. -1

OpenStudy (anonymous):

7. -64 8. 54

OpenStudy (anonymous):

damn you guys beat me to it haha

OpenStudy (anonymous):

:p